OpenStudy (anonymous):

Candace says, "You can write any whole number as the sum of powers of 2 and only use each power once." Is she correct? Try writing 29 as the sum of powers of 2. Then choose two other whole numbers and do the same. Explain your thinking.

OpenStudy (anonymous):

2^4 + 2^3 + 2^2 + 2^0 16 + 8 + 4 + 1 = 29

OpenStudy (anonymous):

Most people forget that 2^0 = 1

OpenStudy (anonymous):

You can take it from here.

OpenStudy (anonymous):

I dont get why you chose those certain exponents. :(

OpenStudy (anonymous):

The problem wanted you to show sums of the power of 2 to equal 29. You could use each power only once so I couldn't do this: 2^3+2^3+2^3+2^2+2^0 8+8+8+4+1= 29 This did't follow the parameters you were given in the problem.

OpenStudy (anonymous):

Let's pick another number: 49. How can you get 49 using the rules outlined in the question?

OpenStudy (anonymous):

the thing is i din't even learn how to add exponents yet

OpenStudy (anonymous):

only multiply and divide

OpenStudy (anonymous):

We're not adding exponents. 2^0 = 1, 2^1 = 2, 2^3 = 8 and so on

OpenStudy (anonymous):

http://www.vaughns-1-pagers.com/computer/powers-of-2.htm check out this table

OpenStudy (anonymous):

oh i get it now

OpenStudy (anonymous):

49 = 2^0 + 2^4 + 2^ 5 49 = 1 + 16 + 32 I chose 2^0 because it equals 1, 49-1=48...then I can break 48 down by powers of 2. I know that 2^5 (2*2*2*2*2)= 32. 48-32 = 16 I know that 2^4 (2*2*2*2)=16

OpenStudy (anonymous):

omg i really understand now, thanks so muchh!! You are my life saver!!! <3